# Multimodal MRI-based Detection of Amyloid Status in Alzheimer’s Disease Continuum

### v1

@@ Abstract
Amyloid- (A) plaques in conjunction with hyperphosphorylated tau proteins in the form of neurofibrillary tangles are the two neuropathological hallmarks of Alzheimer’s disease (AD). In particular, the accumulation of A plaques, as evinced by the A/T/N (amyloid/tau/neurodegeneration) framework, marks the initial stage. Thus, the identification of individuals with A positivity could enable early diagnosis and potentially lead to more effective interventions. Deep learning methods relying mainly on amyloid PET images have been employed to this end. However, PET imaging has some disadvantages, including the need of radiotracers and expensive acquisitions. Hence, in this work, we propose a novel multimodal approach that integrates information from structural, functional, and diffusion MRI data to discriminate A status in the AD continuum. Our method achieved an accuracy of . Furthermore, a post-hoc explainability analysis (guided backpropagation) was performed to retrieve the brain regions that most influenced the model predictions. This analysis identified some key regions that were common across modalities, some of which were well-established AD-discriminative biomarkers and related to A deposition, such as the hippocampus, thalamus, precuneus, and cingulate gyrus. Hence, our study demonstrates the potential viability of MRI-based characterization of A status, paving the way for further research in this domain.
@@ 1 Introduction
The precise identification of A proteins, and thus the classification of patients as A positive or negative, is only possible through amyloid positron emission tomography (PET) scans and cerebrospinal fluid (CSF) tests. Despite their undeniable utility, their use has some negative implications. Specifically, amyloid PET uses ionizing radiation and radioactive tracers, is expensive, and is not widely available ([5, 6]). Additionally, the CSF test requires an invasive lumbar puncture.
@@
In contrast, common, non-invasive MRI techniques can help identifying the effects of A plaques in the brain. In particular, structural MRI (sMRI) has been widely used for AD detection and early-prediction due to its ability to detect GM atrophy and structural changes. Resting-state functional MRI (rs-fMRI) detects changes in blood oxygenation level-dependent (BOLD) signals, which depend on neurovascular coupling, and hence indirectly measures brain neural activity ([7]). Different resting-state networks (RSNs), like the saliency and default-mode (DM) networks, are closely related to AD pathology, showing functional alterations ([8, 9, 10]). Finally, diffusion MRI (dMRI) is an imaging technique that relies on the movement of water molecules to mark WM boundaries enabling the reconstruction of fiber trajectories and the study of brain structural connections in AD ([4]). Additionally, dMRI data are used to extract microstructural information ([11]). These two modalities are generally used to map whole-brain connectomes, with rs-fMRI describing the functional connectivity (FC) between regions (usually in terms of correlations) and dMRI the structural connectivity (SC) (most often relying on the number of WM fibers).
@@
Currently, different studies focusing on A classification tasks (positive vs negative) in AD research have employed PET images and related features with deep learning (DL) models, while only a few works used MRI data. [12] considered a slice-level approach for the identification of the A status in conjunction with a 2D-CNN for feature extraction and classification. In the same manner, [13] adopted the slice-level approach for the classification of A status employing two different 2D-CNN (ResNet-50 and ResNet-152). These two works were able to reach a high level of accuracy (around ) using a slice-level method. On the other hand, [6] considered three different well-known 3D-CNNs for detecting A positivity relying on 3D Florbetaben brain PET images. [14] developed a particular CNN composed of different submodules for analyzing the 3D FDG-PET images, converting them into slices following the three different axes. [15] faced the problem of A classification relying on both early and delay-phase FBB PET images and tested their models considering both single and fused modalities. Using FDG and Amyloid PET-derived images, they were able to reach competitive performance in this task, with accuracies around . More recently, [16] employed FDG-PET-derived features in order to predict the A status in the AD continuum. To this end, they tested eight different feature selection methods and eight different classifiers. LASSO in conjunction with the Gaussian Naive Bayes (GNB) model performed better with respect to the others, achieving an AUC of 0.924. Regarding MRI-based approaches for A detection, [17] employed sMRI images in conjunction with a 3D-CNN, while [18] used an SDF-based convolutional network to analyze the hippocampus region. Using sMRI-derived images and features, they were able to reach accuracies around . Due to the heterogeneous factors that lead to AD, in recent years, many studies have focused on multimodal DL models due to their ability to integrate information of different nature and to outperform single-modality methods ([19, 17, 20]). Moreover, among the different DL models, graph neural and convolutional networks (GNNs/GCNs) are becoming popular in neuroscience due to their perfect fit to functional and structural brain networks. [21] proposed a benchmark for analyzing fMRI and dMRI networks through GNNs, testing different messages passing, node features, and pooling operations, while Wee and colleagues developed a GCN to study the cortical thickness ([22]).
@@
Although DL models can achieve high performance in different tasks, they do not easily provide interpretable output for what they have learned, which is particularly problematic in clinical and biomedical domains. To address this issue, eXplainable Artificial Intelligence (XAI) methods have been developed allowing to identify the contributions of input features to final predictions potentially highlighting crucial physiological information for AD ([23, 24, 25, 26]).
@@
In this study, we present a multimodal and explainable DL framework for the classification of A positive and negative status in the AD continuum. Our framework included structural and functional connectomes derived from dMRI and rs-fMRI, respectively, along with sMRI-derived GM 3D volumes in order to investigate complementary aspects as well as their relations. This was complemented with an extensive post-hoc XAI analysis to uncover the input features that most influenced the final predictions.
2 Methods
@@ 2.1 Dataset
In this work, our dataset was initially composed of preprocessed sMRI images (from subjects) from the ADNI , , , and GO phases, out of which passed quality control (QC) (from subjects). For rs-fMRI, preprocessed images were considered (from individuals) from the ADNI , , and GO phases, out of which passed QC (from individuals). For dMRI, preprocessed images (from subjects) from ADNI were considered, out of which (from subjects) passed QC. Additional information about QC is detailed in the next paragraph. The images of the first available timepoint from only subjects belonging to control (CN), significant memory concern (SMC), early MCI (EMCI), late MCI (LMCI), and AD clinical classes, that had all three modalities, and available A status were included. Lumbar puncture to retrieve CSF samples was performed using the procedures described on the ADNI website, and subjects were labeled as A positive or negative based on the A protein levels reported by the CSF test. Similarly to [5], a cutoff of was used to define the A status (i.e., for positivity). After that, A negative CN, SMC, and EMCI (, , and subjects, respectively) were considered as ’control class’, while A positive EMCI, LMCI, and AD (, , and patients, respectively) formed the ’patient class’, resulting in a final cohort of individuals. Table 1 shows the demographic information of the final cohort.
@@
The MRI images for the considered cohort were collected as follows: i) T1-weighted sMRI: TR/TE=shortest, TI= ms, FOV= mm2, mm isotropic resolution, slices=; ii) rs-fMRI: TR/TE= ms, FOV= mm3, mm isotropic resolution, volumes in almost all subjects, with minimal variations (e.g., ) in a small subset; iii) single-shell dMRI: TR/TE= ms, FOV= mm3, mm isotropic resolution, b= and s/mm2.
@@ 2.3 Framework architecture
The proposed framework is shown in Figure 1. In detail, the DL architecture used for the classification of A status (positive/negative) has two modules: i) a feature reduction module that actuates feature reduction using three different Neural Networks (NNs) to transform the input data into corresponding latent representations; and ii) a data fusion & classification module that concatenates the latent representations of each modality and uses a Multilayer Perceptron (MLP) that takes the fused latent features as input for the final classification. Finally, a post-hoc explainability analysis was performed on the correctly classified subjects to obtain feature contributions for the classification task. In the following paragraphs, the two modules are further described.
@@
Three different NNs form the feature reduction module. Each NN extracts a latent representation of one modality resulting in a latent vector of features. The 3D sMRI volumes were analyzed using a 3D-CNN with six convolutional layers and three max-pooling layers, completed with four dense layers. Conversely, both FNC and SC graphs were analyzed using two different GCNs. These models update the representation of the node aggregating neighbor information, iteratively in each layer, through the message-passing scheme. The rs-fMRI channel was analyzed using a GCN with two graph convolutional layers proposed by [30] that were followed by a dense layer. This type of convolutional layer is a powerful operator that integrates both high- and low-level structures along with their relationships into a single graph ([30]). The dMRI channel was analyzed with a different GCN with two Chebyshev spectral graph convolutional layers proposed by [31] and then followed by a dense layer. Spectral convolutional layers are high-performance layers that have been used effectively in irregular graphs ([32]). These two architectures were chosen on an empirical basis as those leading to the best unimodal classification accuracy.
@@ 2.4 Training scheme and evaluation
The model was trained with stratified -fold cross-validation on the entire cohort to investigate its generalizability across individuals, and a hyperparameter search was performed empirically. The mini-batch strategy (with subjects per batch) was adopted. The Adam optimizer (learning rate: ) was used to update the entire multimodal architecture. L regularization (weight decay: ) was applied to reduce overfitting. Weighted cross entropy was used as the loss function. The model was trained for epochs.
@@
For comparison, we also tested the unimodal models with the corresponding MRI data, where the architectures were the same of the different branches of the multimodal framework (i.e., 3D-CNN for sMRI, and GCN for both rs-fMRI and dMRI).
@@ 3.2 Contribution maps and Statistical analysis
The attribution maps were extracted for the correctly classified A positive subjects, and in particular, we focused on the average A positive attribution map to identify the most important features. To evaluate the sMRI GBP contribution maps, the Harvard-Oxford ([29]) and the probabilistic cerebellar ([35]) atlases from FSL were employed to define different ROIs , including cortical, subcortical, and cerebellum regions. The sum of GBP attributions inside each ROI was calculated for the sMRI and weighted to account for the volume of each specific region. Conversely, the GBP attribution of each node for both rs-fMRI and dMRI was extracted directly from the two GCNs. Then, considering the average map for each modality, the percentage of explanation for each region/node was computed over the total contribution within and between each modality. The top ROIs (sMRI) and nodes (rs-fMRI and dMRI) with the highest percentage of GBP contribution were selected for further investigation.
Subsequent statistical analyses were performed on the original data for all correctly classified subjects. For the sMRI, the mean values of the top ROIs resulting from the XAI analysis were extracted from the input GM volumes and used as features for the statistical analysis. For the rs-fMRI and dMRI, graph-based measures were first derived from the full connectivity matrices in order to have a summary measure per node, and then we considered only the top nodes for both rs-fMRI and dMRI for the statistical analysis. In particular, the node strength was computed for rs-fMRI and FNC matrices. This is defined as the sum of the weights of the edges connected with a given node, where in absolute terms, higher values mean more important nodes. Betweenness centrality was calculated for SC matrices, representing the fraction of all shortest paths in the SC matrix that contains the node under analysis, where the shortest path is the shortest sequence of nodes between node i and node j. As the sparsity of the SC matrix could limit the interpretation of the node strength results for the dMRI data, we preferred to rely on a centrality measure for this analysis. In this case, nodes that belong to more paths likely play a pivotal role in the propagation of the information inside the network. These two metrics were computed using the Brain Connectivity Toolbox (BCT) ([36]) in Matlab.
Mann-Whitney tests were then used to compare the values of all these features for the top ROIs/nodes between A positive and negative correctly classified patients. Finally, FDR correction for multiple comparisons was applied to the p-values obtained from the statistical tests.
@@ 4.1 Classification performance
The single networks that composed the multimodal framework along with the corresponding input data were also tested in the same classification task. Table 2 shows the performance comparisons for the multimodal method and the three unimodal models. Results highlight how the multimodal framework is able to outperform the unimodal models in terms of accuracy, recall, and F1 score in the same classification task.
@@ 4.2 GBP-based attribution maps
The evaluation set in the fold with the highest evaluation accuracy was used for the post-hoc analysis. Importantly, in the analyzed fold, the evaluation subjects were not used for training.
Figure 2A shows the sMRI GBP attribution map for the mean A positive patient overlaid to the MNI template. Qualitatively, mainly subcortical regions (e.g., hippocampus, thalamus) are relevant to the final classification along with a few cortical areas.
@@
Finally, Figure 2C shows the most important dMRI nodes (ROIs). Subcortical regions (thalamus, cerebellum) along with areas in the frontal, temporal, and parietal lobes are identified.

### v2

# Multimodal MRI Accurately Identifies Amyloid Status in Unbalanced Cohorts in Alzheimer’s Disease Continuum

@@ Abstract
Amyloid- (A) plaques in conjunction with hyperphosphorylated tau proteins in the form of neurofibrillary tangles are the two neuropathological hallmarks of Alzheimer’s disease. It is well-known that the identification of individuals with A positivity could enable early diagnosis. In this work, we aim at capturing the A positivity status in an unbalanced cohort enclosing subjects at different disease stages, exploiting the underlying structural and connectivity disease-induced modulations as revealed by structural, functional, and diffusion MRI. Of note, due to the unbalanced cohort, the outcomes may be guided by those factors rather than amyloid accumulation. The partial views provided by each modality are integrated in the model allowing to take full advantage of their complementarity in encoding the effects of the A accumulation, leading to an accuracy of . The specificity of the information brought by each modality is assessed by post-hoc explainability analysis (guided backpropagation), highlighting the underlying structural and functional changes. Noteworthy, well-established biomarker key regions related to A deposition could be identified by all modalities, including the hippocampus, thalamus, precuneus, and cingulate gyrus, witnessing in favor of the reliability of the method as well as its potential in shading light on modality-specific possibly unknown A deposition signatures.
@@ 1 Introduction
In clinical practice, the precise identification of A proteins, and thus the classification of patients as A positive or negative, is performed through amyloid positron emission tomography (PET) scans and cerebrospinal fluid (CSF) tests. Despite their undeniable utility, their use has some inherent limitations. First, limiting to one modality hides informative features that would require other kinds of investigation tools to be revealed. Then, amyloid PET uses ionizing radiation and radioactive tracers, is expensive, and is not widely available ([5, 6]). Additionally, the CSF test requires an invasive lumbar puncture.
@@
A wider perspective on the disease can be obtained by adding views obtained by other imaging modalities. Among these, non-invasive MRI techniques are at the top of the list, allowing to assess structural and functional changes encoding the effect of the amyloid accumulation that can profitably be used as additional biomarkers. Structural MRI (sMRI) has been widely used for AD detection and early prediction due to its ability to detect GM atrophy and structural changes. Resting-state functional MRI (rs-fMRI) detects changes in blood oxygenation level-dependent (BOLD) signals, which depend on neurovascular coupling, and hence indirectly measures brain neural activity ([7]). Different resting-state networks (RSNs), like the saliency and default-mode (DM) networks, have been shown to be altered in AD pathology ([8, 9, 10]). Finally, diffusion MRI (dMRI) is an imaging technique that relies on the movement of water molecules, allowing the assessment of both microstructural ([11]) and structural connectivity (SC) changes ([4]). These two modalities are generally used to map the whole-brain connectomes, with rs-fMRI describing the functional connectivity (FC) between region pairs (usually in terms of correlations) and dMRI the SC (most often relying on the number of WM fibers linking the target regions).
@@
Different studies focusing on A classification tasks (positive vs negative) in AD research have employed PET images and related features with deep learning (DL) models, while only a few works used MRI data. In [12], a slice-level approach for the identification of the A status was considered in conjunction with a 2D-CNN for feature extraction and classification. In the same manner, [13] adopted the slice-level approach for the classification of A status employing two different 2D-CNN (ResNet-50 and ResNet-152), both reaching high accuracy (around ). On the other hand, [6] considered three different well-known 3D-CNNs for detecting A positivity relying on 3D Florbetaben brain PET images. [14] developed a particular CNN composed of different submodules for analyzing the 3D FDG-PET images, converting them into slices following the three different axes. [15] faced the problem of A classification relying on both early and delay-phase FBB PET images and tested their models considering both single and fused modalities. Using FDG and Amyloid PET-derived images, they were able to achieve competitive performance in this task, with accuracies of around . More recently, [16] employed FDG-PET-derived features in order to predict the A status in the AD continuum. To this end, they tested eight different feature selection methods and eight different classifiers. LASSO in conjunction with the Gaussian Naive Bayes (GNB) model performed better with respect to the others, achieving an AUC of . Regarding MRI-based approaches for A detection, [17] employed sMRI images in conjunction with a 3D-CNN, while [18] used an SDF-based convolutional network to analyze the hippocampus region. Using sMRI-derived images and features, they were able to reach accuracies around . Due to the heterogeneous factors that lead to AD, in recent years, many studies have focused on multimodal DL models due to their ability to integrate information of different nature and to outperform single-modality methods ([19, 17, 20]). Recently, graph neural and convolutional networks (GNNs/GCNs) have become popular in neuroscience due to their perfect fit for functional and structural brain networks. [21] proposed a benchmark for analyzing fMRI and dMRI networks through GNNs, testing different messages passing, node features, and pooling operations, while [22] developed a GCN to study the cortical thickness.
@@
Although DL models can achieve high performance in different tasks, they do not easily provide interpretable output for what they have learned, which is particularly problematic in clinical and biomedical domains. To address this issue, eXplainable Artificial Intelligence (XAI) methods have been developed allowing to identify the contributions of input features to final predictions, potentially highlighting crucial information for AD ([23, 24, 25, 26]).
@@
In this study, the overall goal of the proposed approach consists of decrypting the signatures induced by A accumulation in the views provided by s/rs-f/dMRI, taking advantage of their complementarity for capturing the A status while exploiting their interplay for the identification of the regions that played a prominent role in determining the outcome. To instantiate this idea, we present a multimodal and explainable DL framework for the classification of A positive and negative status relying on an unbalanced cohort of individuals spanning the AD continuum. The proposed framework includes structural and functional connectomes derived from dMRI and rs-fMRI, respectively, along with sMRI-derived GM 3D volumes in order to investigate complementary aspects as well as their relations. For the sake of interpretability of the results, an extensive post-hoc XAI analysis was then performed, pointing the spotlight on the input features that most influenced the final outcome.
2 Materials and Methods
@@ 2.1 Dataset
In this work, our dataset was initially composed of preprocessed sMRI images (from subjects) from the ADNI , , , and GO phases, out of which passed quality control (QC) (from subjects). For rs-fMRI, preprocessed images were considered (from individuals) from the ADNI , , and GO phases, out of which passed QC (from individuals). For dMRI, preprocessed images (from subjects) from ADNI were considered, out of which (from subjects) passed QC. Additional information about QC is detailed in the next paragraph. The images of the first available timepoint from only subjects belonging to control (CN), significant memory concern (SMC), early MCI (EMCI), late MCI (LMCI), and AD clinical classes, that had all three modalities, and available A status were included. Lumbar puncture to retrieve CSF samples was performed using the procedures described on the ADNI website, and subjects were labeled as A positive or negative based on the A protein levels reported by the CSF test. Similarly to [5], a cutoff of was used to define the A status (i.e., for positivity).
Aiming at the classification based on the A status as the target outcome, the considered group of individuals was further split gathering the A negative CN, SMC, and EMCI individuals in the NEG class (, , and subjects), and the A positive EMCI, LMCI, and AD in the POS class (, , and subjects), respectively, resulting in an unbalanced data split with respect to the disease stage. The LMCI and AD groups were not included in the A negative class since they could represent different underlying conditions linked with the functional decline (e.g., Lewy Body dementia, Frontotemporal dementia, vascular dementia, TDP-43 pathology). In contrast, CN and SMC individuals with amyloid accumulation are subjects of ongoing debate, with no clear consensus on whether they represent a prodromal stage of AD or individuals at higher risk of developing the disease but still may never develop the disease ([27, 28]).
@@
The MRI images for the considered cohort were collected as follows: i) T1-weighted sMRI: TE/TR=shortest, TI= ms, FOV= mm2, mm isotropic resolution, slices=; ii) rs-fMRI: TE/TR= ms, FOV= mm3, mm isotropic resolution, volumes in almost all subjects, with minimal variations (e.g., ) in a small subset; iii) single-shell dMRI: TE/TR= ms, FOV= mm3, mm isotropic resolution, b= and s/mm2.
@@ 2.3 Framework architecture
The proposed framework is shown in Figure 1. In detail, the DL architecture used for the classification of A status (positive/negative) has two modules: i) a feature reduction module that actuates feature reduction using three different Neural Networks (NNs) to transform the input data into corresponding latent representations; and ii) a data fusion & classification module that concatenates the latent representations of each modality and uses a Multilayer Perceptron (MLP) that takes the fused latent features as input for the final classification. Finally, a post-hoc explainability analysis was performed on the correctly classified subjects to highlight the feature contributions to the classification task. In the following paragraphs, the two modules are further described.
@@
Three different NNs form the feature reduction module. Each NN extracts a latent representation of one modality resulting in a latent vector of features. The 3D sMRI volumes were analyzed using a 3D-CNN with six convolutional layers and three max-pooling layers, completed with four dense layers. Conversely, both FNC and SC graphs were analyzed using two different GCNs. These models update the representation of each node, aggregating neighbor information iteratively in each layer through the message-passing scheme. The rs-fMRI channel was analyzed using a GCN with two graph convolutional layers proposed by [32] that were followed by a dense layer. This type of convolutional layer is a powerful operator that integrates both high- and low-level structures along with their relationships into a single graph ([32]). The dMRI channel was analyzed with a different GCN with two Chebyshev spectral graph convolutional layers proposed in [33] followed by a dense layer. Spectral convolutional layers are high-performance layers that have been used effectively on irregular graphs ([34]). These two architectures were chosen on an empirical basis as those leading to the best unimodal classification accuracy.
@@ 2.4 Training scheme and evaluation
The model was trained with stratified -fold cross validation on the entire cohort to investigate its generalizability across individuals, and a hyperparameter search was performed empirically through a grid search procedure to maximize the average validation accuracy. To this end, different combinations were tested for the hyperparameters (batch size, learning rate, number of epochs, and regularization parameter). Additionally, different numbers of hidden layers for CNN, GCNs, and classifier were considered, also changing the number of channels for the convolutional layers and the number of neurons in the dense layers. The mini-batch strategy (with subjects per batch) was finally adopted. The Adam optimizer (learning rate: ) was used to update the entire multimodal architecture. L regularization (weight decay: ) was applied to reduce overfitting. Weighted cross entropy was used as the loss function. The model was trained for epochs.
@@
For the sake of comparison, we also tested the unimodal models with the corresponding MRI data, where the architectures were the same as the different branches of the multimodal framework (i.e., 3D-CNN for sMRI, and GCN for both rs-fMRI and dMRI).
@@ 3.2 Contribution maps and Statistical analysis
The attribution maps were extracted for the correctly classified A positive subjects. The average A positive subject attribution map was derived for identifying the most important features. To evaluate the sMRI GBP contribution maps, the Harvard-Oxford ([31]) and the probabilistic cerebellar ([37]) atlases from FSL were employed to define different ROIs, including cortical, subcortical, and cerebellum regions. The sum of GBP attributions inside each ROI was calculated for the sMRI and weighted to account for the volume of each specific region. Conversely, the GBP attribution of each node for both rs-fMRI and dMRI was extracted directly from the two GCNs. Then, considering the average map for each modality, the percentage of explanation for each region/node was computed over the total contribution within and across modalities. The top ROIs (sMRI) and nodes (rs-fMRI and dMRI) with the highest percentage of GBP contribution were selected for further investigation.
Subsequent statistical analyses were performed on the original data for all correctly classified subjects. For the sMRI, the mean values of the top ROIs resulting from the XAI analysis were extracted from the input GM volumes and used as features for the statistical analysis. For the rs-fMRI and dMRI, graph-based measures were first derived from the full connectivity matrices in order to have a summary measure per node, and then only the top nodes were retained for both rs-fMRI and dMRI for statistical analysis. In particular, the node strength was computed for the FNC matrices. This is defined as the sum of the weights of the edges connected with a given node, where in absolute terms, higher values mean more important nodes. Betweenness centrality was calculated for SC matrices, representing the fraction of all shortest paths in the SC matrix that contains the node under analysis, where the shortest path is the shortest sequence of nodes between node i and node j. As the sparsity of the SC matrix could limit the interpretation of the node strength results for the dMRI data, we preferred to rely on a centrality measure for this analysis. In this case, nodes that belong to more paths likely play a pivotal role in the propagation of the information inside the network. These two metrics were computed using the Brain Connectivity Toolbox (BCT) ([38]) in Matlab.
Mann-Whitney tests were then used to compare the values of all these features for the top ROIs/nodes between A positive and negative correctly classified individuals. Finally, FDR correction for multiple comparisons was applied.
@@ 4.1 Classification performance
The single networks that composed the multimodal framework along with the corresponding input data were also tested in the same classification task. Table 2 shows the performance comparisons for the multimodal and the three unimodal models. Results highlight how the multimodal pipeline is able to outperform the unimodal models in terms of accuracy, recall, and F1 score in the same classification task.
@@ 4.2 GBP-based attribution maps
The evaluation set in the fold with the highest evaluation accuracy was used for post-hoc analysis. Importantly, in the analyzed fold, the evaluation and training sets were disjoint.
Figure 2A shows the sMRI GBP attribution map for the A mean positive subject, overlaid to the MNI template. Qualitatively, mainly subcortical regions (e.g., hippocampus, thalamus) were relevant to the final classification along with a few cortical areas.
@@
Finally, Figure 2C shows the most important dMRI nodes (ROIs). Subcortical regions (thalamus, cerebellum) along with areas in the frontal, temporal, and parietal lobes were identified.